determine the mean,median and mode of the following sets of number
a)8,11,14,13,14,9,15

Answer by lawdej(58)

mean=(8+11+14+13+14+9+15)/7=84/7=12
mode=14
median
first arrange in order
8,9,11,13,14,14,15 so 13 is the median